# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

### 10 Mar 2026 v1

@@ 6.3 Efficiency-Driven Applications in Federated and Distributed Learning
These distributed settings significantly reshape how model combination can be performed. Model merging assumes particular significance in such contexts where computational resources are geographically dispersed and data privacy constraints preclude centralized training. Federated learning represents the canonical application scenario, wherein multiple client devices or institutions independently fine-tune local model copies on private datasets, subsequently aggregating these models at a central server without transmitting raw training data. The foundational Federated Averaging (FedAvg) algorithm (McMahan et al., 2017) implements precisely this paradigm, performing weighted averaging of client model parameters according to local dataset sizes: $\theta_{\text{global}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\theta_{k}$, where $n_{k}$ denotes the number of samples at client $k$ and $n$ represents the total sample count across all participants. This formulation directly instantiates the uniform averaging principles discussed in Section 3, connecting federated learning directly to weight-space interpolation theory.
@@
The extension of federated learning to large language models introduces substantial challenges absent in smaller-scale settings. Communication bandwidth constraints become particularly acute when synchronizing models with billions of parameters across potentially unreliable network connections. To address this limitation, researchers have developed communication-efficient aggregation protocols that transmit only sparse parameter updates or low-rank approximations of weight differences rather than complete model states (McMahan et al., 2017). These approaches leverage the sparsification principles underlying TIES-Merging and DARE, transmitting trimmed task vectors that capture essential fine-tuning modifications while dramatically reducing bandwidth requirements. The mathematical formulation adapts naturally: clients compute local task vectors $\tau_{k}=\theta_{k}-\theta_{\text{base}}$, apply sparsification operators to retain only high-magnitude parameters, and transmit these compressed representations for server-side aggregation.
@@
The branch-train-merge paradigm offers an alternative distributed training strategy particularly suited for large-scale model development. Rather than continuously synchronizing parameters during training, this approach allows independent branches to diverge substantially before periodic merging operations consolidate progress. This strategy proves especially effective when training specialized experts on distinct data domains, as the reduced synchronization frequency enables more efficient utilization of heterogeneous computational resources. The theoretical viability of branch-train-merge depends critically on maintaining mode connectivity between diverging branches, a property that shared pretrained initialization generally ensures for moderate fine-tuning dur.[truncated]
@@
Deployment optimization through model consolidation represents another efficiency-driven application with significant practical implications. Organizations frequently maintain multiple specialized models fine-tuned for distinct tasks or domains, incurring substantial inference costs from running parallel model instances. Deployment efficiency is critically important for edge computing scenarios where computational resources are severely constrained. Model merging enables consolidating multiple specialized variants into unified systems that serve diverse requirements through a single inference pathway, dramatically reducing memory footprint and operational costs. This consolidation proves particularly valuable for edge deployment scenarios where resource limitations prohibit maintaining separate model instances for different functionalities.
@@
Scalability considerations present fundamental challenges for merging operations across large participant populations. The computational complexity of pairwise alignment procedures, required when clients lack shared initialization, scales quadratically with participant count, making naive approaches intractable at scale. Hierarchical aggregation strategies address this limitation through recursive merging of participant subgroups, reducing effective complexity while potentially sacrificing solution quality. Furthermore, the statistical heterogeneity characteristic of federated settings, wherein local data distributions exhibit substantial divergence, can compromise mode connectivity assumptions underlying successful merging, necessitating adaptive aggregation strategies that account for inter-client distribution shifts. While these distributed and efficiency-driven applications address computational and privacy constraints, model merging also serves a complementary purpose in reconciling specialized domain expertise with general-purpose capabilities.
@@ 6.4 Domain Specialization While Preserving General Capabilities
Domain specialization exemplifies this reconciliation challenge, presenting a fundamental tension in the development of large language models: while foundation models exhibit broad competence across diverse tasks, achieving expert-level performance in specialized domains such as healthcare, legal, scientific, or cybersecurity applications typically requires focused fine-tuning that risks degrading general-purpose capabilities. Model merging has emerged as a principled approach for resolving this tension, enabling the integration of domain-specific expertise while preserving the broad knowledge encoded in pretrained foundations. The mathematical formulation underlying this capability balance can be expressed as $\theta_{\text{merged}}=(1-\alpha)\theta_{\text{base}}+\alpha\theta_{\text{domain}}$, where the interpolation coefficient $\alpha$ governs the trade-off between domain expertise and general versatility.
@@
Healthcare applications exemplify the critical importance of this capability preservation challenge. As comprehensively surveyed in (Yang et al., 2024b), medical language models must navigate the dual requirements of clinical accuracy and broad conversational competence. Fine-tuning foundation models on clinical corpora, including electronic health records, medical literature, and diagnostic guidelines, produces models with enhanced biomedical reasoning but frequently compromised performance on general language understanding tasks. Model merging addresses this limitation by combining clinically fine-tuned variants with their general-purpose ancestors, producing unified systems that maintain diagnostic capability while preserving the linguistic fluency essential for patient communication. The interpolation coefficient in healthcare merging scenarios typically favors domain expertise ($\alpha\in[0.6,0.8]$) for clinical decision support applications while emphasizing general capabilities ($\alpha\in[0.3,0.5]$) for patient-facing conversational interfaces.
Cybersecurity applications present analogous challenges with distinct technical characteristics. Multiple specialized capabilities are required for security applications, including vulnerability detection, malware analysis, and threat intelligence processing. Similarly, language models can be adapted for both offensive and defensive security tasks. Merging strategies for cybersecurity domains must balance technical security knowledge with general reasoning capabilities, as effective security analysis frequently requires synthesizing domain-specific threat models with broader contextual understanding. Unlike healthcare applications where domain boundaries are relatively well-defined, cybersecurity merging often involves combining models specialized for distinct security subtasks, code vulnerability detection, network traffic analysis, and social engineering detection, into unified security assistants.
@@
Legal and regulatory compliance domains introduce additional complexity through the requirement for jurisdiction-specific knowledge preservation during merging. Models fine-tuned on legal corpora from different jurisdictions may encode conflicting regulatory frameworks, necessitating careful management of task vector interactions during the merging process. The task vector formulation provides natural mechanisms for such management: representing jurisdiction-specific adaptations as separate vectors enables controlled combination through weighted aggregation that respects regulatory boundaries. Recent advances in data-free merging demonstrate that sophisticated guidance mechanisms can achieve superior performance without requiring access to original training distributions, a property particularly valuable for legal applications where training data may be subject to confidentiality constraints.
@@
Scientific and technical domains benefit from hierarchical merging strategies that progressively specialize from general scientific reasoning to field-specific expertise. A foundation model might first be merged with a broadly scientific variant before incorporating narrower specializations in chemistry, physics, or biology. This hierarchical approach leverages the compositional properties of task arithmetic while managing the increased interference potential arising from multiple simultaneous specializations. The interpolation coefficient optimization for scientific merging often employs validation sets spanning both general scientific literacy and domain-specific benchmarks, enabling automated discovery of Pareto-optimal coefficient configurations that maximize the joint performance frontier. Across all these domain specialization scenarios, the effectiveness of merging strategies ultimately depends on systematic assessment methodologies capable of quantifying both capability retention and cross-domain interference.
@@ 6.5 Evaluation Methodologies and Benchmarking for Merged Model Assessment
Given the critical dependence of domain-specific merging strategies on systematic assessment, the rigorous evaluation of merged models presents unique methodological challenges that extend beyond standard model assessment paradigms, requiring frameworks capable of capturing capability retention, multi-task performance, and interference effects simultaneously. Unlike evaluation of single-task fine-tuned models, merged model assessment must quantify both the preservation of individual source model capabilities and the emergent properties arising from their combination, necessitating evaluation protocols that probe along multiple orthogonal dimensions.
@@
Standardized benchmarking for merged models typically employs composite evaluation suites that aggregate performance across the constituent tasks represented in the source models. The evaluation protocol computes task-specific performance metrics $P_{i}^{\text{merged}}$ for each task $i$ in the set of source tasks $\mathcal{T}$, comparing against the performance of individual fine-tuned models $P_{i}^{\text{ft}}$ to compute retention ratios $R_{i}=P_{i}^{\text{merged}}/P_{i}^{\text{ft}}$. The aggregate retention metric, often computed as the geometric mean across tasks $R_{\text{geo}}=(\prod_{i\in\mathcal{T}}R_{i})^{1/|\mathcal{T}|}$, provides a single scalar summary that penalizes severe degradation on any individual task while remaining interpretable. Empirical evaluations demonstrate that advanced merging methods such as TIES-Merging and task arithmetic approaches can achieve retention ratios exceeding 0.95 on certain NLP classification tasks; however, this strong retention is most consistently observed when merging models of similar scale (typically 7B parameters or fewer), when source tasks exhibit semantic relatedness, and when employing interference-aware merging strategies. Performance degradation becomes more pronounced for knowledge-intensive tasks, generative benchmarks, and scenarios involving dissimilar task combinations, where retention ratios may fall to 0.80–0.90 or lower.
@@
Concrete benchmark suites have become essential for standardized merged model evaluation. For language models, commonly employed benchmarks include MMLU (Hendrycks et al., 2021) for broad knowledge assessment, HellaSwag (Zellers et al., 2019) for commonsense reasoning, WinoGrande (Sakaguchi et al., 2020) for coreference resolution, ARC (Clark et al., 2018) for scientific reasoning, GSM8K (Cobbe et al., 2021) for mathematical problem-solving, and TruthfulQA (Lin et al., 2022) for factual accuracy and hallucination detection. These benchmarks serve complementary evaluation purposes: MMLU and ARC probe knowledge retention across diverse domains, HellaSwag and WinoGrande assess reasoning capabilities that may benefit from cross-task transfer, while GSM8K and TruthfulQA evaluate specialized competencies that are particularly susceptible to interference during merging. For code generation capabilities, HumanEval (Chen et al., 2021) serves as a standard benchmark.
@@
Interference detection methodologies focus on identifying performance degradation patterns that emerge specifically from the merging process rather than from limitations of individual source models. A principled interference metric computes the pairwise performance difference $\Delta_{ij}=P_{i}^{\text{merged}(i,j)}-P_{i}^{\text{merged}(i)}$ comparing merged model performance on task $i$ when combined with task $j$ versus in isolation. Systematic negative values indicate destructive interference, while positive values suggest beneficial cross-task transfer. The interference matrix $\mathbf{I}$ with entries $I_{ij}=\Delta_{ij}$ provides a view of task interactions, enabling identify problematic task combinations and guide selective merging decisions. Recent work on representation analysis provides complementary tools for understanding how merged representations differ from their source models at intermediate layers.
@@
Cross-task generalization assessment extends evaluation beyond source task performance to probe whether merged models exhibit capabilities that transcend their constituent specializations. This evaluation dimension addresses the hypothesis that successful merging produces models with compositional generalization, the ability to combine knowledge from different source tasks to solve novel problems. Evaluation protocols for compositional generalization employ held-out tasks that require simultaneous application of multiple specialized capabilities, measuring whether the merged model outperforms any individual source model on these composite benchmarks. Research on weight averaging (Izmailov et al., 2018) demonstrated that weight averaging can yield improved out-of-distribution generalization, motivating subsequent investigation into the generalization properties of more sophisticated merging schemes.
@@
Beyond task-specific accuracy, comprehensive merged model evaluation increasingly incorporates calibration assessment, out-of-distribution robustness testing, and computational efficiency comparisons. Calibration evaluation measures whether merged model confidence scores accurately reflect empirical correctness probabilities, as merging may distort the calibration properties of source models even when accuracy is preserved. Expected calibration error (ECE) and reliability diagrams provide standard tools for this assessment. Out-of-distribution robustness testing evaluates merged model performance on inputs drawn from distributions distinct from source task training data, using benchmarks designed to probe distributional shift sensitivity. Computational efficiency comparisons quantify inference latency, memory footprint, and throughput characteristics relative to ensemble baselines and individual source models, contextualizing accuracy gains against resource requirements.
@@
Community evaluation has converged on leaderboard-based methodologies that standardize comparison across research groups. Standardized evaluation frameworks (Alcover-Couso et al., 2024) have emerged for reproducible LLM evaluation, supporting consistent benchmark execution across diverse model architectures. The Open LLM Leaderboard hosted by Hugging Face aggregates performance across standardized benchmark suites (He et al., 2025), enabling direct comparison of merged models against conventionally trained baselines and facilitating community-driven progress tracking. Model hub infrastructure on platforms such as Hugging Face additionally supports merged model sharing, enabling reproduction and extension of published results. Recent investigations into evaluation methodology highlight the importance of statistical rigor in merged model assessment, advocating for confidence intervals and significance testing when comparing merging methods given the variance inherent in benchmark evaluation. The adoption of multi-objective evaluation frameworks that simultaneously assess helpfulness, harmlessness, and honesty dimensions reflects the understanding that merged model quality cannot be captured by a single metric, especially in safety-critical deployments.
FusionBench (Tang et al., 2024b) addresses this gap by providing a unified evaluation suite covering 25+ merging methods across language tasks, enabling consistent cross-method comparison. It introduces task retention rate (TRR), defined as $\text{TRR}=\frac{1}{K}\sum_{i=1}^{K}\frac{\text{Perf}_{\text{merged},i}}{\text{Perf}_{\text{source},i}}$, as a standardized metric quantifying how much of each source model’s capability survives the merge. On LLM benchmarks, TRR provides actionable guidance for practitioners comparing simple averaging against advanced interference-aware methods.
@@
Despite these advances, significant evaluation gaps persist in the merged model literature. Standardized benchmarks specifically designed to measure interference effects remain lacking, forcing researchers to rely on indirect comparison methodologies that may obscure subtle capability degradation. Evaluation protocols for emergent capabilities, skills present in merged models but absent from all source models—remain poorly developed, limiting understanding of when and why beneficial emergence occurs. These persistent challenges notwithstanding, the collective progress across merging methodologies, theoretical understanding, and evaluation practices provides a foundation for synthesizing the current state of the field and identifying the most promising directions for future research.
Table 8: Comparison of application scenarios and practical approaches in model merging.

### 30 Mar 2026 v2

@@ 6.3 Efficiency-Driven Applications in Federated and Distributed Learning
These distributed settings impose distinctive constraints on model combination. Model merging assumes particular significance in such contexts where computational resources are geographically dispersed and data privacy constraints preclude centralized training. Federated learning represents the canonical application scenario, wherein multiple client devices or institutions independently fine-tune local model copies on private datasets, subsequently aggregating these models at a central server without transmitting raw training data. The foundational Federated Averaging (FedAvg) algorithm (McMahan et al., 2017) implements precisely this approach by performing weighted averaging of client model parameters according to local dataset sizes. $\theta_{\text{global}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\theta_{k}$, where $n_{k}$ denotes the number of samples at client $k$ and $n$ represents the total sample count across all participants. This formulation directly instantiates the uniform averaging principles discussed in Section 3, connecting federated learning directly to weight-space interpolation theory.
@@
The extension of federated learning to large language models introduces unique challenges absent in smaller-scale settings. Communication bandwidth constraints become particularly acute when synchronizing models with billions of parameters across potentially unreliable network connections. To mitigate this limitation, researchers have developed communication-efficient aggregation protocols that transmit only sparse parameter updates or low-rank approximations of weight differences rather than complete model states (McMahan et al., 2017). These approaches exploit the sparsification principles underlying TIES-Merging and DARE, transmitting trimmed task vectors that capture essential fine-tuning modifications while dramatically reducing bandwidth requirements. The mathematical formulation adapts naturally. Clients compute local task vectors $\tau_{k}=\theta_{k}-\theta_{\text{pre}}$, apply sparsification operators to retain only high-magnitude parameters, and transmit these compressed representations for server-side aggregation.
@@
The branch-train-merge strategy offers an alternative distributed training approach particularly suited for large-scale model development. Rather than continuously synchronizing parameters during training, this approach allows independent branches to diverge considerably before periodic merging operations consolidate progress. This strategy proves especially effective when training specialized experts on distinct data domains, as the reduced synchronization frequency permits more efficient use of heterogeneous computational resources. The theoretical viability of branch-train-merge depends critically on maintaining mode connectivity between diverging branches, a property that shared pretrained initialization generally ensures for moderate fine-tuning durations.
@@
Deployment optimization through model consolidation offers direct cost savings. Organizations frequently maintain multiple specialized models for distinct tasks or domains, incurring costs from running parallel instances. Model merging consolidates these variants into a single system serving diverse requirements through one inference pathway, reducing memory footprint proportionally to the number of consolidated models. This consolidation proves particularly valuable for edge deployment scenarios where resource limitations prohibit maintaining separate model instances.
@@
Scalability considerations present major challenges for merging operations across large participant populations. The computational complexity of pairwise alignment procedures, required when clients lack shared initialization, scales quadratically with participant count, making naive approaches intractable at scale. Hierarchical aggregation strategies alleviate this limitation through recursive merging of participant subgroups, reducing effective complexity while potentially sacrificing solution quality. Mediator (Lai et al., 2025) proposes a memory-efficient LLM merging approach that reduces parameter conflicts through mediation, making practical merging at scale feasible with less computational overhead. Moreover, the statistical heterogeneity characteristic of federated settings, wherein local data distributions exhibit marked divergence, can compromise mode connectivity assumptions underlying successful merging, necessitating adaptive aggregation strategies that account for inter-client distribution shifts. While these distributed and efficiency-driven applications handle computational and privacy constraints, model merging also serves a complementary purpose in reconciling specialized domain expertise with general-purpose capabilities.
@@ 6.4 Domain Specialization While Preserving General Capabilities
Domain specialization exemplifies this reconciliation challenge; while foundation models exhibit broad competence across diverse tasks, achieving expert-level performance in specialized domains such as healthcare, legal, scientific, or cybersecurity applications typically requires focused fine-tuning that risks degrading general-purpose capabilities. Model merging offers a principled means to resolve this tension by integrating domain-specific expertise while preserving the broad knowledge encoded in pretrained foundations. The mathematical formulation underlying this capability balance can be expressed as $\theta_{\text{merged}}=(1-\alpha)\theta_{\text{pre}}+\alpha\theta_{\text{domain}}$, where the interpolation coefficient $\alpha$ governs the trade-off between domain expertise and general versatility.
@@
Healthcare applications exemplify the critical importance of this capability preservation challenge. As thoroughly surveyed in (Yang et al., 2024b), medical language models must navigate the dual requirements of clinical accuracy and broad conversational competence. Fine-tuning foundation models on clinical corpora, including electronic health records, medical literature, and diagnostic guidelines, produces models with enhanced biomedical reasoning but frequently compromised performance on general language understanding tasks. Model merging remedies this limitation by combining clinically fine-tuned variants with their general-purpose ancestors, producing unified systems that maintain diagnostic capability while preserving the linguistic fluency essential for patient communication. Timilsina et al. (2025) provide a systematic evaluation of six parameter-space merging techniques applied to architecturally compatible medical LLMs derived from Mistral-7B, finding that Task Arithmetic achieves 45.80% accuracy on MedQA and that simple averaging provides a robust and computationally efficient baseline for distributed medical AI in resource-constrained environments. The interpolation coefficient in healthcare merging scenarios is typically tuned to favor domain expertise for clinical decision support applications while emphasizing general capabilities for patient-facing conversational interfaces.
Cybersecurity represents a promising but largely unexplored application domain. Security applications require multiple specialized capabilities, including vulnerability detection, malware analysis, and threat intelligence processing, each of which can benefit from domain-specific fine-tuning. Merging strategies could in principle balance technical security knowledge with general reasoning, as effective security analysis frequently requires synthesizing domain-specific threat models with broader contextual understanding. However, no published model merging studies in this domain have appeared, leaving the effectiveness of current methods for security-specific knowledge composition as an open empirical question.
@@
Legal and regulatory compliance domains present a conceptually appealing but under-explored scenario. Models fine-tuned on legal corpora from different jurisdictions may encode conflicting regulatory frameworks, necessitating careful management of task vector interactions. The task vector formulation naturally supports such management, as jurisdiction-specific adaptations represented as separate vectors permit controlled combination through weighted aggregation. Data-free merging methods may be particularly relevant where training data is subject to confidentiality constraints, though no published studies have evaluated model merging for legal domain specialization to date.
@@
Scientific and technical domains benefit from hierarchical merging strategies that progressively specialize from general scientific reasoning to field-specific expertise. A foundation model might first be merged with a broadly scientific variant before incorporating narrower specializations in chemistry, physics, or biology. This hierarchical approach exploits the compositional properties of task arithmetic while managing the increased interference potential arising from multiple simultaneous specializations. Siriwardhana et al. (2024) provide a detailed evaluation of domain adaptation through continual pre-training and model merging for Llama3-70B-Instruct, confirming the practical viability of merging-based domain adaptation at scale. Hirano and Imajo (2024) demonstrate the construction of instruction-tuned LLMs for finance without instruction data using continual pretraining and model merging, showcasing how merging can bypass the need for domain-specific instruction datasets. The interpolation coefficient optimization for scientific merging often uses validation sets spanning both general scientific literacy and domain-specific benchmarks, enabling automated discovery of Pareto-optimal coefficient configurations that maximize the joint performance frontier. Across all these domain specialization scenarios, the effectiveness of merging strategies ultimately depends on systematic assessment methodologies capable of quantifying both capability retention and cross-domain interference.
@@ 6.5 Evaluation Methodologies and Benchmarking for Merged Model Assessment
Evaluating merged models presents unique methodological challenges that extend beyond standard model assessment. Unlike evaluation of single-task fine-tuned models, merged model assessment must quantify both the preservation of individual source model capabilities and the emergent properties arising from combination, requiring evaluation protocols that probe along multiple orthogonal dimensions.
@@
Standardized benchmarking for merged models typically uses composite evaluation suites that aggregate performance across the constituent tasks represented in the source models. The evaluation protocol computes task-specific performance metrics $P_{i}^{\text{merged}}$ for each task $i$ in the set of source tasks $\mathcal{T}$, comparing against the performance of individual fine-tuned models $P_{i}^{\text{ft}}$ to compute retention ratios $R_{i}=P_{i}^{\text{merged}}/P_{i}^{\text{ft}}$. The aggregate retention metric, often computed as the geometric mean across tasks $R_{\text{geo}}=(\prod_{i\in\mathcal{T}}R_{i})^{1/|\mathcal{T}|}$, provides a single scalar summary that penalizes severe degradation on any individual task while remaining interpretable. Empirical evaluations show that advanced merging methods such as TIES-Merging and task arithmetic approaches can produce high retention ratios on many NLP tasks; however, strong retention is most consistently observed when merging models of similar scale, when source tasks exhibit semantic relatedness, and when employing interference-aware merging strategies. Performance degradation becomes more pronounced for knowledge-intensive tasks, generative benchmarks, and scenarios involving dissimilar task combinations.
@@
Concrete benchmark suites have become essential for standardized merged model evaluation. For language models, commonly used benchmarks include MMLU (Hendrycks et al., 2021) for broad knowledge assessment, HellaSwag (Zellers et al., 2019) for commonsense reasoning, WinoGrande (Sakaguchi et al., 2020) for coreference resolution, ARC (Clark et al., 2018) for scientific reasoning, GSM8K (Cobbe et al., 2021) for mathematical problem-solving, and TruthfulQA (Lin et al., 2022) for factual accuracy and hallucination detection. These benchmarks serve complementary evaluation purposes. MMLU and ARC probe knowledge retention across diverse domains, HellaSwag and WinoGrande assess reasoning capabilities that may benefit from cross-task transfer, while GSM8K and TruthfulQA evaluate specialized competencies that are particularly susceptible to interference during merging. For code generation capabilities, HumanEval (Chen et al., 2021) serves as a standard benchmark.
@@
Interference detection methodologies focus on identifying performance degradation patterns that emerge specifically from the merging process rather than from limitations of individual source models. A principled interference metric computes the pairwise performance difference $\Delta_{ij}=P_{i}^{\text{merged}(i,j)}-P_{i}^{\text{merged}(i)}$ comparing merged model performance on task $i$ when combined with task $j$ versus in isolation. Systematic negative values indicate destructive interference, while positive values suggest beneficial cross-task transfer. The interference matrix $\mathbf{I}$ with entries $I_{ij}=\Delta_{ij}$ provides a view of task interactions, enabling identification of problematic task combinations and guiding selective merging decisions.
@@
Cross-task generalization assessment extends evaluation beyond source task performance to probe whether merged models exhibit capabilities that transcend their constituent specializations. This evaluation dimension probes the hypothesis that successful merging produces models with compositional generalization, the ability to combine knowledge from different source tasks to solve novel problems. Evaluation protocols for compositional generalization use held-out tasks that require simultaneous application of multiple specialized capabilities, measuring whether the merged model outperforms any individual source model on these composite benchmarks. Research on weight averaging (Izmailov et al., 2018) showed that weight averaging can yield improved out-of-distribution generalization, motivating subsequent investigation into the generalization properties of more advanced merging schemes.
@@
Beyond task-specific accuracy, merged model evaluation increasingly incorporates calibration assessment, out-of-distribution robustness testing, and computational efficiency comparisons. Calibration evaluation measures whether merged model confidence scores accurately reflect empirical correctness probabilities, as merging may distort the calibration properties of source models even when accuracy is preserved. Expected calibration error (ECE) and reliability diagrams provide standard tools for this assessment. Out-of-distribution robustness testing evaluates merged model performance on inputs drawn from distributions distinct from source task training data, using benchmarks designed to probe distributional shift sensitivity. Computational efficiency comparisons quantify inference latency, memory footprint, and throughput characteristics relative to ensemble baselines and individual source models, contextualizing accuracy gains against resource requirements.
@@
Community evaluation has converged on leaderboard-based methodologies that standardize comparison across research groups. The Open LLM Leaderboard hosted by Hugging Face aggregates performance across standardized benchmark suites, permitting direct comparison of merged models against conventionally trained baselines and enabling community-driven progress tracking. Model hub infrastructure on platforms such as Hugging Face additionally supports merged model sharing, allowing reproduction and extension of published results. Recent investigations into evaluation methodology highlight the importance of statistical rigor in merged model assessment, advocating for confidence intervals and significance testing when comparing merging methods given the variance inherent in benchmark evaluation. The adoption of multi-objective evaluation frameworks that simultaneously assess helpfulness, harmlessness, and honesty dimensions reflects the understanding that merged model quality cannot be captured by a single metric, especially in safety-critical deployments.
Dedicated merging benchmarks have advanced evaluation rigor. FusionBench (Tang et al., 2024c) provides a unified evaluation suite covering 25+ merging methods across language tasks, enabling consistent cross-method comparison. MergeBench (He et al., 2025b) specifically targets domain-specialized LLM merging evaluation. Tam et al. (2024b) present a realistic evaluation framework focusing on compositional generalization, revealing that many reported gains may not transfer to more challenging settings. Hitit et al. (2025) conduct a large-scale evaluation of six merging methods across four open-weight LLMs, finding that Task Arithmetic is the only approach that reliably yields performance gains, while more recent methods frequently produce marked performance drops, cautioning against assuming advances on smaller models generalize to LLMs. Cao et al. (2026b) provide both empirical and theoretical analysis of task-level model-merging collapse, identifying conditions under which merged models catastrophically fail on specific tasks despite maintaining aggregate performance. A useful standardized metric is the task retention rate (TRR), defined as $\text{TRR}=\frac{1}{K}\sum_{i=1}^{K}\frac{\text{Perf}_{\text{merged},i}}{\text{Perf}_{\text{source},i}}$, which quantifies how much of each source model’s capability survives the merge (Yadav et al., 2025b).
@@
Despite these advances, notable evaluation gaps persist in the merged model literature. Standardized benchmarks specifically designed to measure interference effects remain lacking, forcing researchers to rely on indirect comparison methodologies that may obscure subtle capability degradation. Evaluation protocols for emergent capabilities, skills present in merged models but absent from all source models, remain poorly developed, limiting understanding of when and why beneficial emergence occurs. These persistent challenges notwithstanding, the collective progress across merging methodologies, theoretical understanding, and evaluation practices provides a foundation for synthesizing the current state of the field and identifying the most promising directions for future research.
Table 8: Application scenarios for model merging in LLM development. Each scenario is categorized by its primary motivation (capability augmentation, alignment/safety, or efficiency), key merging approach, distinctive features, and known limitations. Year indicates when the approach was first applied in the LLM context.
@@
The diverse application scenarios examined in this section show that model merging has matured from a theoretical curiosity into a practical technique with tangible impact across capability augmentation, safety alignment, federated learning, and domain specialization, while the evaluation methodologies discussed provide essential frameworks for rigorously assessing merged model quality. Emerging application directions continue to expand the scope of merging across multiple frontiers, though the maturity and evidence base vary considerably across domains.
Reasoning and chain-of-thought.
Model merging has shown particular promise for reasoning enhancement. A unifying theme is the long-to-short distillation strategy: models trained with extended chain-of-thought are merged with more efficient counterparts to transfer reasoning ability without the inference cost. Examples include efficient long-to-short LLM reasoning (Wu et al., 2025a), fine-grained reasoning injection (Huang et al., 2026a), model interpolation for efficient reasoning (Wu et al., 2025b), reasoning pattern alignment (Zhong et al., 2026), merging long chain-of-thought models with domain-specific models (Yang et al., 2026b), and tunable reasoning via the thinking spectrum (Lan et al., 2025). Notably, Kimi k1.5 (Team et al., 2025) illustrates industrial-scale deployment of this approach, using model merging to transfer long-chain reasoning into efficient short-chain models. The key tradeoff across these methods is between reasoning depth (preserved through larger merging coefficients for the reasoning model) and inference efficiency (favored by smaller coefficients).
Continual and lifelong learning.
Merging offers natural mechanisms for continual adaptation by framing knowledge accumulation as sequential model combination rather than catastrophic retraining. The central challenge is forgetting: each merge risks overwriting previously acquired knowledge. Methods addressing this include aligned model merging for vision-language models (Sokar et al., 2025), streaming LLM updates via activation-guided rotations (Yao et al., 2026), null-space filtering for data-free continual merging (Qiu et al., 2025b), K-Merge for online continual adaptation (Shenaj et al., 2025), and RECALL for catastrophic-forgetting alleviation (Wang et al., 2025a). The diversity of these approaches, ranging from null-space projection to activation-guided rotation, reflects the absence of a single dominant solution, suggesting that the interaction between merging and continual learning remains an open research area.
Multimodal and cross-modal.
Merging enables integration across modalities, addressing the practical need to combine separately trained vision, language, and action modules without joint retraining. Representative works include ES-Merging for biological MLLM merging (Lee et al., 2026), linear model merging for multimodal data mixture optimization (Berasi et al., 2026), OptMerge for unifying multimodal capabilities (Wei et al., 2025b), transferring textual preferences to VLMs (Li et al., 2025a), RobustMerge for parameter-efficient MLLM merging (Zeng et al., 2025), Graft for domain knowledge integration (Dai et al., 2025), and test-time merging for zero-shot medical imaging (Imam et al., 2025). A common finding is that cross-modal merging is more sensitive to representation alignment than unimodal merging, since different modalities occupy fundamentally different regions of parameter space.
Safety and intellectual property.
Beyond the alignment applications discussed in Section 6, emerging safety-relevant directions highlight a dual-use concern: the same merging operations that enable beneficial knowledge composition can also introduce adversarial risks. H3Fusion (Tekin et al., 2024) addresses multi-objective alignment by fusing models optimized for helpfulness, harmlessness, and honesty, split-unlearn-merge (Kadhe et al., 2024) addresses selective knowledge removal, mitigating social biases through unlearning (Dige et al., 2024) targets fairness, alignment-preserving Fisher-guided merging (Roy et al., 2025) provides safety guarantees, and defending unauthorized merging (Chen et al., 2025e) and copyright protection (Abad et al., 2024) address intellectual property concerns. Notably, unlearning via model merging (Xu et al., 2025a) shows that TIES-Merging can balance over-forgetting and under-forgetting by merging complementary bias models, demonstrating that merging’s interference properties can be harnessed constructively for safety objectives.
Cross-lingual and agentic applications.
Cross-lingual transfer via layer swapping (Bandarkar et al., 2025) and catastrophic forgetting mitigation in language transfer (Alexandrov et al., 2024) extend merging to multilingual settings, while language-specific model merging (Dmonte et al., 2026) shows that independently training per-language models and merging them via task arithmetic can reduce training and maintenance costs compared to joint multilingual training. Agentic applications represent a newer frontier: role-conditioned neuron transplantation (Feng et al., 2026), behavior knowledge merge for reinforced agents (Yuan et al., 2026), LLM steering via persona vectors (Pai et al., 2025), and GTR-Turbo (Wei et al., 2025a), which merges checkpoint weights from RL training to create a free teacher model for self-distillation, collectively suggest that merging can serve as a lightweight mechanism for composing behavioral capabilities in autonomous systems.
Scaling and meta-analysis.
Foundational studies on scaling reveal both promise and limitations: model merging scaling laws (Wang et al., 2025c) characterize how merging benefits evolve with model size, while empirical analysis of what matters for merging at scale (Yadav et al., 2025b) identifies key factors. Model-GLUE (Zhao et al., 2024a) and Merge to Learn (Morrison et al., 2024) show democratized scaling through merging. Multi-objective merging (Li et al., 2024a; Zhou et al., 2024a; Xie et al., 2025b; a) addresses the realistic scenario where practitioners must balance multiple competing objectives, while knowledge fusion through semantic alignment (Gu et al., 2025) and weight evolution (Du et al., 2024b) explore alternative combination strategies.